Ratios for XYZ

The following information for 2021 and 2020 is presented for XYZ

 




December   31





Assets


2021

2020

Current   assets:









Cash

$

42,000

$     54,000

Accounts   receivable


580,000

445,000

Inventory


5,010,000

4,950,000

Prepaid expenses


84,000


79,000

Total   current assets


5,716,000

5,528,000

Building   and equipment, net


1,097,000


1,095,000

Total   assets

$

6,813,000

$6,623,000

Liabilities and Stockholders’ Equity

















Current liabilities:









Accounts   payable

$

605,000

$  628,000

Bank   loan payable


679,000

625,000

Other accrued payables


215,000


315,000

Total   current liabilities


1,499,000

1,568,000

Long-term   debt


1,729,000

1,791,000







Total   liabilities


3,228,000

3,359,000

Stockholders’ equity:









Common   stock


1,307,000

1,307,000

Retained earnings


2,278,000


1,957,000

Total   stockholders’ equity


3,585,000


3,264,000

Total   liabilities and stockholders’ equity

$

6,813,000

$6,623,000










 

There were 100,000 shares of common stock outstanding throughout both 2020 and 2021.

 

Additional information follows:

 



2021



2020











Market price per share at   the end of year

$


134

$

110

Net income for the year


815,000


639,000

Cost of goods sold for   the year


2,900,000


2,700,000

Net sales for the year


5,568,000


5,253,000










 

1.            Calculate the 1) current ratio, 2) acid-test ratio, and the 3) debt-to-equity ratio for 2020 and 2021. Calculate to three significant digits.

 

2.            The company intends to apply for a loan. What concerns might the loan officer have about lending to the company? 
Answer #1

SOLUTION :


Ratios :


1.


Current ratio (CR ) 

=  Current Assets, CA / Current Liabilities, CL


So,


CR (2020) = 5528000/1568000 = 3.526 (ANSWER)


CR(2021) = 5716000/1499000 = 3.813 9ANSWER).


2.


Acid test ratio (ATR)

= (CA - Inventories) / CL


So,


ATR (2020) = (5528000 - 4950000) / 1568000 = 0.369 (ANSWER)


ATR (2021) = (5716000 - 5010000) / 1499000 = 0.471(ANSWER)


3.


Debt-equity ratio (DE) 

=Total liabilities / Equity


So,


DE (2020) = 3359000 / 3264000 = 1.029 (ANSWER)


DE (2021) = 3228000 / 3585000 = 0.900 (ANSWER).



Concerns for lending :


Current ratio is greater than 3 in both 2020 and 2921 . It is good. However, ATR is much lower than 1.0 in both the years and seems to be deteriorating. This indicates poor liquidity condition of the company. In case of liquidation, current liabilities could not be neutralised by cash like assets. There is too much of inventories. As regards DE ratio, It is around 1.0 and improving over 2020 in 2021. It is reasonably alright.


Hence, low ATR and too much of inventories are the concern in lending loan to the company. There is risk that company may face cash crunch and may fail to service the loan repay it. 



answered by: Tulsiram Garg